# MoGe: Unlocking Accurate Monocular Geometry Estimation for Open-Domain Images with Optimal Training Supervision

### v2

@@ 4 Experiments
LeReS 17.0 76.0 37.3 11.1 17.1 75.8 18.5 72.1 14.7 75.9 38.6 30.6 31.9 39.6 27.5 46.7 25.3 53.5 3.94
DUSt3R 5.56 97.1 21.9 63.6 10.7 90.5 6.17 95.4 4.55 99.3 34.8 50.4 21.3 70.3 12.3 87.0 14.7 81.7 2.69
UniDepth 5.36 98.3 10.7 98.0 18.5 77.5 5.27 97.4 6.57 99.6 33.0 49.0 11.3 90.3 12.3 91.1 12.9 87.7 2.18
Ours 4.91 98.4 8.62 96.3 4.60 98.9 4.62 97.1 2.60 100 22.3 69.4 12.3 90.4 6.50 94.7 8.32 93.1 1.12
@@
LeReS 9.55 91.4 29.0 43.8 14.7 79.6 11.0 88.7 8.89 95.3 29.7 55.5 29.3 46.8 14.9 80.3 18.4 72.7 3.94
DUSt3R 4.49 97.4 18.0 66.7 7.27 94.9 5.02 96.0 3.08 99.5 30.3 56.7 19.6 71.4 8.86 88.9 12.1 83.9 2.94
UniDepth 3.96 98.4 7.94 97.3 12.2 89.6 4.62 98.0 2.98 99.8 28.5 58.5 10.2 90.6 8.46 91.1 9.86 90.4 1.88
Ours 3.73 98.3 7.69 95.0 3.60 99.0 3.60 97.3 1.16 100 16.9 77.8 10.4 91.5 4.31 96.6 6.43 94.4 1.25
@@
LeReS - - 8.62 91.5 15.7 84.4 - 18.7 85.8 1.65 99.0 16.2 85.2 12.2 89.2 3.70
DUSt3R - - 5.88 94.4 9.43 91.3 - 12.7 88.3 1.71 98.9 10.1 91.3 7.96 92.9 2.30
UniDepth - - 7.45 92.7 10.2 90.2 - 17.0 86.9 1.64 98.8 13.2 87.8 9.89 91.3 3.00
Ours - - 2.86 97.9 7.02 94.1 - 9.73 92.1 1.08 99.5 5.96 95.2 5.33 95.8 1.00
@@
LeReS 12.1 82.5 19.3 64.3 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.1 18.1 69.9 18.5 69.5 7.31
ZoeDepth 5.65 96.3 7.23 92.2 10.4 87.3 7.48 93.2 3.23 99.9 27.4 61.7 17.0 73.0 11.2 85.4 11.2 86.1 5.50
DUSt3R 4.43 97.1 7.71 90.9 6.07 95.6 4.96 95.9 3.27 99.5 31.1 57.2 18.5 73.5 8.84 88.8 10.6 87.3 5.00
Metric3D V2 4.74 97.3 3.96 98.6 3.92 98.5 4.20 97.7 2.47 99.9 20.7 70.0 7.36 94.6 3.27 98.5 6.32 94.4 2.07
UniDepth 3.88 98.3 3.73 98.7 5.68 97.0 4.76 97.4 4.17 99.7 28.3 58.9 9.98 90.7 6.79 92.9 8.41 91.7 3.00
DA V1 4.81 97.5 5.52 95.8 8.80 92.3 5.54 95.8 5.48 99.2 28.1 56.7 13.1 81.7 10.1 87.8 10.2 88.4 5.67
DA V2 5.09 97.2 7.28 93.6 6.14 95.5 4.33 97.9 4.40 99.3 22.9 65.3 14.6 78.2 7.89 90.3 9.08 89.7 4.06
Ours 3.48 98.4 4.27 97.8 3.38 98.9 3.45 97.1 1.49 100 19.4 73.4 9.11 90.6 4.85 94.8 6.17 93.9 1.62
@@
MiDaS V3.1 4.62 98.1 6.30 94.8 5.80 96.8 4.75 97.4 1.88 100 21.4 73.0 14.4 82.7 6.06 94.9 8.15 92.2 3.69
DA V1 4.24 98.4 5.40 97.0 4.70 98.2 4.20 97.6 1.54 100 20.4 77.5 12.6 87.0 5.72 95.7 7.36 93.9 2.31
DA V2 4.20 98.2 5.61 96.8 4.73 98.0 3.48 98.5 1.25 100 21.4 72.8 13.0 86.6 5.33 96.1 7.38 93.4 2.50
Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.94 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.06
Table 2: Quantitative results for depth map estimation. Gray numbers denote models trained on respective benchmarks and thus excluded from ranking. An extensive comparison is found in the suppl. materials.
@@
Depth+camera 11.3 88.7 9.20 90.4 10.1 91.2 8.97 90.1 7.29 92.6 8.22 92.1
Opt. scale-inv. 10.5 89.7 8.50 91.4 9.09 92.0 8.28 90.9 6.72 93.3 7.91 92.6
Opt. affine-inv. 10.1 90.3 8.07 91.9 7.84 93.2 7.92 91.2 6.29 93.7 7.43 93.2
Med. affine-inv. 11.5 88.5 9.36 89.7 10.2 90.9 9.11 89.7 7.38 92.4 8.53 92.0
Full w/o truncation 10.1 90.4 8.13 91.4 6.94 94.0 7.94 91.4 6.31 93.5 7.45 93.1
Full w/o $\mathcal{L}_{S}$ 10.1 90.2 8.10 91.8 7.51 93.5 7.94 91.2 6.31 93.6 7.47 93.2
Full 10.0 90.5 8.00 91.8 7.07 94.0 7.84 91.4 6.21 93.8 7.31 93.3
Table 4: Quantitative ablation study results. All experiments are conducted with a ViT-Base backbone for the encoder.
@@
4.1 Quantitative Evaluations
We assess the zero-shot performance of our model and compare it to several state-of-the-art methods for monocular point map estimation [71, 42, 59] and depth map estimation [67, 68, 6, 26, 5, 20, 29]. To ensure a fair comparison, all methods utilize ViT-Large [15] as backbone, except for LeReS (using ResNeXt101 [65]) as well as GeoWizard and Marigold (using StableDiffusion V1 [46]). The scale and shift of the predicted point maps and depth maps are aligned with the ground truth when necessary. We also evaluate the FOV prediction results against several recent methods [28, 77]. As detailed below, our method achieves superior results across various tasks and benchmarks.
@@ 4.1 Quantitative Evaluations
We evaluate the accuracy of point map estimation on eight diverse datasets: NYUv2 [39], KITTI [53], ETH3D [50], iBims-1 [31], Sintel [9], Google Scanned Objects (GSO) [16], DDAD [24], and DIODE [54]. These datasets represent a wide range of domains, including indoor scenes, street views, object scans, and synthetic movies. The raw datasets are processed for reliable evaluation (e.g., sky regions in Sintel dataset and boundary artifacts in the DIODE are removed; see the suppl. material for more preprocessing details). For evaluation metrics, we use the relative point error Relp, i.e., $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}$, and the percentage of inliers $\delta_{1}^{\text{p}}$ with $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}<0.25$.
@@
To further assess local geometry accuracy, we evaluate affine-invariant point maps within regions defined by object segmentation masks, sourced from dataset labels or Segment Anything [30]. Due to the need for high-quality ground truth and scenes with multiple objects, NYUv2, KITTI, and GSO are excluded. Our method achieves a substantial reduction in region-wise Relp, lowering the error from $7.96$ to $5.33$—an approximately $30\%$ improvement over previous methods.
@@ 4.3 Ablation Study
We compare our affine-invariant point map representation with two variants, one scale-invariant point map in camera space, and another that predicts a depth map and camera parameters. For the former, we apply the proposed ROE alignment to solve the scale parameter only. For the latter, we adopt the SI-log loss [18] for the depth map and the MSE loss for the ray direction map, similar to UniDepth [42]. Results are shown in the first three rows of Table 4, where only the global ground-truth supervision is applied in these experiments. Our proposed affine-invariant representation outperforms the others across various tasks. This is attributed to the fact that the other two variants suffer from focal-distance ambiguity issue during training, which hinders their performance.
@@
To evaluate the effectiveness of our optimal alignment, we compare it to a previously-used simple alignment method [43], i.e., the predicted and ground-truth point maps are normalized by their median: $(\mathbf{p}-\mathbf{t}(\mathbf{P}))/s(\mathbf{P})$, where $\mathbf{t}(\mathbf{P})$ equals to $(0,0,\operatorname{median}(\mathbf{Z}))$, and $s(\mathbf{P})$ is computed as $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}-\mathbf{t}(\mathbf{P})\|_{1}$. As shown in the third and fourth rows of Table 4, our proposed alignment method consistently outperforms the median alignment strategy across all tasks by large margins, demonstrating the critical role of our proposed optimal alignment.
We ablate the truncation of residuals in alignment objects, which degrades performance (see the fifth row in Table 4). In practice, adding truncation also improves training stability.
@@
As shown in the six-th row of Table 4, the quantitative results without $\mathcal{L}_{S}$ suffer a significant drop on local geometry metrics. This demonstrates the effectiveness of $\mathcal{L}_{S}$ in enhancing local geometry accuracy and compensating for the insufficient supervision of the global loss due to the ambiguous relative positions between distant scene objects. Fig. 6 (a) further demonstrates the effectiveness of the proposed local geometry loss.
@@ D. Kinect Medium $\sim 10$m Dense
Table 5: Datasets used for training (top) and the tailored loss combinations (bottom).
@@ B.3 Evaluation Protocol
  • Scale-invariant depth map. The scale $a^{*}$ is computed as
@@ Appendix C More Results
Full table of depth estimation results.
@@
LeReS 12.1 82.5 19.3 64.3 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.1 18.1 69.9 18.5 69.5 7.31
ZoeDepth 5.65 96.3 7.23 92.2 10.4 87.3 7.48 93.2 3.23 99.9 27.4 61.7 17.0 73.0 11.2 85.4 11.2 86.1 5.50
DUSt3R 4.43 97.1 7.71 90.9 6.07 95.6 4.96 95.9 3.27 99.5 31.1 57.2 18.5 73.5 8.84 88.8 10.6 87.3 5.00
Metric3D V2 4.74 97.3 3.96 98.6 3.92 98.5 4.20 97.7 2.47 99.9 20.7 70.0 7.36 94.6 3.27 98.5 6.32 94.4 2.07
UniDepth 3.88 98.3 3.73 98.7 5.68 97.0 4.76 97.4 4.17 99.7 28.3 58.9 9.98 90.7 6.79 92.9 8.41 91.7 3.00
DA V1* 4.81 97.5 5.52 95.8 8.80 92.3 5.54 95.8 5.48 99.2 28.1 56.7 13.1 81.7 10.1 87.8 10.2 88.4 5.67
 - metric indoor 4.81 97.5 15.2 74.2 9.42 89.0 5.54 95.8 5.48 99.2 28.3 56.7 24.2 57.5 10.1 87.8 12.9 82.2 -
 - metric outdoor 16.0 72.2 5.52 95.8 8.80 92.3 13.8 78.8 8.60 93.6 28.1 54.8 13.1 81.7 12.9 81.7 13.3 81.4 -
DA V2* 5.09 97.2 7.28 93.6 6.14 95.5 4.33 97.9 4.40 99.3 22.9 65.3 14.6 78.2 7.89 90.3 9.08 89.7 4.06
 - metric indoor 5.09 97.2 7.49 91.3 6.14 95.5 4.33 97.9 4.40 99.3 22.9 65.3 16.5 73.6 7.89 90.3 9.35 88.8 -
 - metric outdoor 15.3 72.1 7.28 93.6 9.33 89.6 10.6 84.9 9.61 92.5 28.6 57.4 14.6 78.2 12.1 83.6 13.4 81.5 -
Ours 3.48 98.4 4.27 97.8 3.38 98.9 3.45 97.1 1.49 100 19.4 73.4 9.11 90.6 4.85 94.8 6.17 93.9 1.62
@@
LeReS 6.25 95.4 8.15 90.6 8.98 90.7 6.71 94.6 4.03 99.4 24.0 64.8 16.1 75.9 9.93 88.4 10.5 87.5 8.81
ZoeDepth 4.79 97.2 5.51 95.3 7.29 94.2 5.88 95.7 2.54 99.9 21.8 69.2 14.1 80.3 7.72 91.0 8.70 90.3 7.25
DUSt3R 3.76 97.8 7.21 91.8 4.96 96.4 3.92 96.6 2.56 99.6 25.4 64.1 16.8 76.4 6.55 92.9 8.89 89.5 6.69
Metric3D V2 4.00 97.6 3.48 98.5 3.54 98.4 3.25 98.3 2.12 99.4 24.4 71.6 7.10 94.8 2.73 98.8 6.33 94.7 3.50
UniDepth 3.42 98.6 3.54 98.8 4.93 97.5 3.74 98.2 2.47 99.9 24.8 64.1 9.36 91.0 4.88 96.2 7.14 93.0 3.62
Marigold 4.63 97.3 7.29 93.8 6.08 96.3 4.35 97.2 2.78 99.9 21.2 75.0 14.6 80.5 6.34 94.3 8.41 91.8 5.81
GeoWizard* 4.69 97.4 8.14 92.5 6.90 94.0 4.50 97.1 2.00 99.9 17.8 76.2 16.5 75.7 7.03 92.7 8.44 90.7 6.56
- indoor 4.69 97.4 8.14 92.5 6.90 93.9 4.50 97.1 2.00 99.9 17.8 76.0 16.5 75.7 7.03 92.7 8.44 90.7 -
- outdoor 4.83 97.3 8.40 92.2 7.18 93.7 4.59 97.0 2.03 99.9 17.8 76.2 17.1 74.4 7.22 92.5 8.64 90.4 -
- object 4.70 97.4 8.21 92.4 6.92 94.0 4.52 97.0 2.05 99.9 17.9 76.0 16.8 75.1 7.13 92.6 8.52 90.5 -
DA V1* 3.86 98.3 4.94 96.6 6.26 96.5 4.24 97.3 1.98 100 20.1 71.8 11.2 86.2 6.50 93.4 7.55 92.8 4.50
 - metric indoor 3.86 98.3 9.78 87.2 6.35 95.0 4.24 97.3 1.98 100 20.1 71.8 16.9 74.2 6.50 93.0 8.72 89.6 -
 - metric outdoor 7.72 93.8 4.94 96.6 6.26 96.5 7.01 94.2 2.78 99.8 20.7 70.0 11.2 86.2 6.93 93.4 8.44 91.3 -
DA V2* 4.22 97.8 6.81 94.2 4.66 97.2 3.45 98.3 1.45 100 17.2 76.5 14.2 82.0 5.32 94.9 7.16 92.6 3.19
 - metric indoor 4.22 97.8 6.95 92.7 4.66 97.2 3.45 98.3 1.45 100 17.2 76.5 14.2 80.0 5.32 94.9 7.18 92.2 -
 - metric outdoor 8.69 90.9 6.81 94.2 7.26 93.5 6.81 93.5 2.28 100 22.5 67.0 13.2 82.0 8.00 91.2 9.45 89.0 -
Ours 2.97 98.6 3.95 98.0 2.71 99.2 2.73 97.9 0.96 100 13.1 83.1 8.33 92.2 3.13 97.6 4.73 95.8 1.56

### v3

@@ 4 Experiments
LeReS 16.9 76.0 31.6 28.4 17.1 75.8 18.5 72.2 14.7 76.0 38.6 30.6 32.0 39.4 27.6 46.4 24.6 55.6 3.94
DUSt3R 5.53 97.1 15.2 87.9 10.7 90.6 6.18 95.4 4.54 99.3 34.8 50.3 21.4 70.1 12.4 86.7 13.8 84.7 2.75
UniDepth 5.33 98.4 5.96 98.5 18.5 77.6 5.29 97.4 6.58 99.6 33.0 48.9 11.4 90.2 12.3 91.0 12.3 87.7 2.09
Ours 4.86 98.4 5.47 97.4 4.58 98.9 4.63 97.1 2.58 100 22.3 69.5 12.3 90.3 6.58 94.5 7.91 93.3 1.22
@@
LeReS 9.51 91.4 26.1 49.1 14.7 79.6 11.0 88.6 8.91 95.2 29.7 55.5 29.4 46.7 15.1 80.1 18.1 73.3 3.94
DUSt3R 4.45 97.4 12.7 83.3 7.27 95.0 5.04 96.0 3.07 99.6 30.3 56.6 19.7 71.2 8.97 88.7 11.4 86.0 2.94
UniDepth 3.93 98.4 4.29 98.6 12.2 89.6 4.65 98.0 2.99 99.8 28.5 58.4 10.3 90.5 8.56 90.9 9.43 90.5 1.81
Ours 3.68 98.3 4.86 97.2 3.57 99.0 3.61 97.3 1.14 100 16.8 77.8 10.5 91.4 4.37 96.4 6.07 94.7 1.31
@@
LeReS - - - - 9.32 91.9 8.57 93.2 - - 13.3 84.8 10.7 88.9 11.6 88.2 10.7 89.4 3.80
DUSt3R - - - - 6.05 94.8 5.44 95.9 - - 11.8 87.0 9.24 90.8 7.32 93.1 7.97 92.3 2.30
UniDepth - - - - 8.61 92.6 5.92 96.0 - - 13.4 84.3 8.18 92.0 9.95 90.0 9.21 91.0 2.90
Ours - - - - 3.21 98.1 4.16 96.8 - - 8.63 92.7 6.74 94.3 4.78 96.3 5.50 95.6 1.00
@@
LeReS 12.1 82.6 19.2 64.8 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.0 18.2 69.6 18.5 69.5 7.31
ZoeDepth 5.62 96.3 7.27 91.9 10.4 87.3 7.45 93.2 3.23 99.9 27.4 61.8 17.0 72.8 11.3 85.2 11.2 86.1 5.50
DUSt3R 4.40 97.1 7.81 90.6 6.04 95.7 4.98 95.8 3.27 99.5 31.1 57.2 18.6 73.3 8.91 88.8 10.6 87.2 5.00
Metric3D V2 4.69 97.4 4.00 98.5 3.84 98.5 4.23 97.7 2.46 99.9 20.7 69.8 7.41 94.6 3.29 98.4 6.33 94.3 2.07
UniDepth 3.86 98.4 3.73 98.6 5.67 97.0 4.79 97.4 4.18 99.7 28.3 58.8 10.1 90.5 6.83 92.8 8.43 91.6 3.00
DA V1 4.77 97.5 5.61 95.6 9.41 88.9 5.53 95.8 5.49 99.3 28.3 56.7 13.2 81.5 10.3 87.5 10.3 87.9 5.67
DA V2 5.03 97.3 7.23 93.7 6.12 95.5 4.32 97.9 4.38 99.3 23.0 65.2 14.7 78.0 7.95 90.0 9.09 89.6 4.06
Ours 3.44 98.4 4.25 97.8 3.36 98.9 3.46 97.0 1.47 100 19.3 73.4 9.17 90.5 4.89 94.7 6.17 93.8 1.62
@@
MiDaS V3.1 4.58 98.1 6.25 94.7 5.77 96.8 4.73 97.4 1.86 100 21.3 73.1 14.5 82.6 6.05 94.9 8.13 92.2 3.69
DA V1 4.20 98.4 5.40 97.0 4.68 98.2 4.18 97.6 1.54 100 20.1 77.6 12.7 86.9 5.69 95.7 7.31 93.9 2.31
DA V2 4.14 98.3 5.61 96.7 4.71 97.9 3.47 98.5 1.24 100 21.4 72.8 13.1 86.4 5.29 96.1 7.37 93.3 2.56
Ours 3.38 98.6 4.05 98.1 3.11 98.9 3.23 98.0 0.96 100 18.4 79.5 8.99 91.5 3.98 97.2 5.76 95.2 1.06
Table 2: Quantitative results for depth map estimation. Gray numbers denote models trained on respective benchmarks and thus excluded from ranking. A more extensive comparison can be found in the suppl. materials.
@@
SI-Log depth 11.2 88.7 9.09 90.6 9.19 91.2 8.94 90.1 7.27 92.6 8.23 92.1
Affine-inv. depth 29.9 51.4 29.0 52.7 12.2 86.0 28.9 52.7 6.18 93.9 15.9 76.6
ROE scale-inv. 10.3 89.8 8.34 91.6 8.59 91.9 8.27 90.9 6.73 93.2 7.90 92.6
L2 affine-inv. 13.5 84.2 10.3 88.2 9.48 91.0 11.1 85.7 8.03 91.2 9.37 90.5
Med. affine-inv. 10.9 89.0 8.97 90.7 9.44 90.7 9.10 89.8 7.50 92.4 8.74 91.8
ROE affine-inv. 9.84 90.3 7.88 92.1 7.62 93.3 7.91 91.2 6.29 93.7 7.43 93.2
Full w/o trunc. 9.81 90.5 7.91 91.7 7.12 93.8 7.92 91.3 6.31 93.5 7.45 93.1
Full w/o $\mathcal{L}_{S}$ 9.98 90.3 7.94 92.1 7.47 93.4 7.94 91.2 6.30 93.6 7.47 93.2
Full 9.78 90.6 7.83 92.1 7.16 93.8 7.82 91.3 6.20 93.7 7.30 93.3
Table 4: Quantitative ablation study results. All experiments are conducted with a ViT-Base encoder. The first six rows are trained with global loss only.
@@
4.1 Quantitative Evaluations.
We assess the zero-shot performance of our model and compare it to several state-of-the-art methods for monocular point map estimation [71, 42, 59] and depth map estimation [67, 68, 6, 26, 5, 20, 29]. To ensure a fair comparison, all methods utilize ViT-Large [15] as backbone, except for LeReS (using ResNeXt101 [65]) as well as GeoWizard and Marigold (using StableDiffusion V1 [46]). The scale and shift of the predicted point maps and depth maps are aligned with the ground truth when necessary. We also evaluate the FOV prediction results against several recent methods [28, 77]. As detailed below, our method achieves superior results across various tasks and benchmarks. [^1] In this revision, we corrected evaluation numbers due to a minor error in the evaluation data. The overall rankings and conclusions are not affected. All results are reproducible using the released code.
@@
We evaluate the accuracy of point map estimation on eight diverse datasets: NYUv2 [39], KITTI [53], ETH3D [50], iBims-1 [31], Sintel [9], Google Scanned Objects (GSO) [16], DDAD [24], and DIODE [54]. These datasets represent a wide range of domains, including indoor scenes, street views, object scans, and synthetic movies. The raw datasets are processed for reliable evaluation (e.g., sky regions in Sintel dataset and boundary artifacts in the DIODE are removed; see the suppl. material for more preprocessing details). For evaluation metrics, we use the relative point error Relp, i.e., $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\|\mathbf{p}\|_{2}$, and the percentage of inliers $\delta_{1}^{\text{p}}$ with $\|\hat{\mathbf{p}}-\mathbf{p}\|_{2}/\min(\|\mathbf{p}\|_{2},\|\hat{\mathbf{p}}\|_{2})<0.25$.
@@
To further assess local geometry accuracy, we evaluate affine-invariant point maps within regions defined by object segmentation masks, sourced from dataset labels or Segment Anything [30]. Due to the need for high-quality ground truth and scenes with multiple objects, NYUv2, KITTI, and GSO are excluded. Our method achieves a substantial reduction in region-wise Relp, lowering the error from $7.97$ to $5.50$—an approximately $30\%$ improvement over previous methods.
@@ 4.3 Ablation Study
We compare our affine-invariant point map representation with three variants: (1) a scale-invariant depth map and a dense ray map for camera parameters, (2) an affine-invariant depth map with a shift scalar and a dense ray map, and (3) a scale-invariant point map in camera space. The first variant uses SI-log loss [18] for the depth map and MSE for the ray map, follows UniDepth [42]. The second employs ROE alignment with $L_{1}$ loss for the affine depth, $L_{1}$ loss for shift, and MSE for the ray map. The third applies the ROE alignment to solve the scale parameter only. Table 4 (row 1-3) shows that our affine-invariant representation outperforms the others across various tasks. Notably, while the affine-invariant depth representation slightly improves specific metrics, all others drop dramatically. Our representation excels by effectively resolving focal-distance ambiguity, a key issue that hampers effectively training of the other three variants.
@@
To assess the effectiveness of our optimal alignment, we compare it to least-squares alignment for $L_{2}$ loss and a previously-used median-based alignment [43] for $L_{1}$ loss. The former has an analytic solution, while the latter normalizes predicted and ground-truth point maps by their median: $(\mathbf{p}-\mathbf{t}(\mathbf{P}))/s(\mathbf{P})$, where $\mathbf{t}(\mathbf{P})=(0,0,\operatorname{median}(\mathbf{Z}))$, and $s(\mathbf{P})=\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}-\mathbf{t}(\mathbf{P})\|_{1}$. As shown in the Table 4 (row 4-6), our proposed alignment method consistently outperforms the other two strategies across all tasks by large margins, highlighting its critical role. We also ablate the truncation of alignment objective residuals, which degrades performance (Table 4 row 7).
@@
As shown in Table 4 (row 8), the quantitative results without $\mathcal{L}_{S}$ suffer a significant drop on local geometry metrics. This demonstrates the effectiveness of $\mathcal{L}_{S}$ in enhancing local geometry accuracy and compensating for the insufficient supervision of the global loss due to the ambiguous relative positions between distant scene objects. Fig. 6 (a) further demonstrates the effectiveness of the proposed local geometry loss.
@@ D. Kinect Medium $\sim 10$m Dense
Table 5: Datasets used for training and tailored loss combination.
@@ B.3 Evaluation Protocol
  • Scale-invariant depth map, the scale $a^{*}$ is computed as
@@ Appendix C More Results
Full table of depth estimation results
@@
LeReS 12.1 82.6 19.2 64.8 14.2 78.4 14.0 78.8 13.6 77.9 30.5 52.1 26.5 52.0 18.2 69.6 18.5 69.5 7.31
ZoeDepth 5.62 96.3 7.27 91.9 10.4 87.3 7.45 93.2 3.23 99.9 27.4 61.8 17.0 72.8 11.3 85.2 11.2 86.1 5.50
DUSt3R 4.40 97.1 7.81 90.6 6.04 95.7 4.98 95.8 3.27 99.5 31.1 57.2 18.6 73.3 8.91 88.8 10.6 87.2 5.00
Metric3D V2 4.69 97.4 4.00 98.5 3.84 98.5 4.23 97.7 2.46 99.9 20.7 69.8 7.41 94.6 3.29 98.4 6.33 94.3 2.07
UniDepth 3.86 98.4 3.73 98.6 5.67 97.0 4.79 97.4 4.18 99.7 28.3 58.8 10.1 90.5 6.83 92.8 8.43 91.6 3.00
DA V1 4.77 97.5 5.61 95.6 9.41 88.9 5.53 95.8 5.49 99.3 28.3 56.7 13.2 81.5 10.3 87.5 10.3 87.9 5.67
  -metric indoor 4.77 97.5 15.4 73.6 9.41 88.9 5.53 95.8 5.49 99.3 28.3 56.7 24.2 57.4 10.3 87.5 12.9 82.1 -
  -metric outdoor 15.9 72.3 5.61 95.6 8.77 92.4 13.8 78.8 8.59 93.6 28.1 54.8 13.2 81.5 13.0 81.4 13.4 81.3 -
DA V2 5.03 97.3 7.23 93.7 6.12 95.5 4.32 97.9 4.38 99.3 23.0 65.2 14.7 78.0 7.95 90.0 9.09 89.6 4.06
  -metric indoor 5.03 97.3 7.61 90.9 6.12 95.5 4.32 97.9 4.38 99.3 23.0 65.2 16.6 73.4 7.95 90.0 9.38 88.7 -
  -metric outdoor 15.3 72.3 7.23 93.7 9.30 89.6 10.6 84.9 9.62 92.5 28.6 57.3 14.7 78.0 12.2 83.2 13.4 81.4 -
Ours 3.44 98.4 4.25 97.8 3.36 98.9 3.46 97.0 1.47 100 19.3 73.4 9.17 90.5 4.89 94.7 6.17 93.8 1.62
@@
LeReS 6.21 95.4 8.28 90.3 8.95 90.8 6.68 94.5 4.03 99.4 24.0 64.8 16.2 75.8 9.99 88.1 10.5 87.4 8.81
ZoeDepth 4.76 97.3 5.59 95.1 7.27 94.2 5.85 95.7 2.54 99.9 21.8 69.2 14.2 80.1 7.80 90.9 8.73 90.3 7.33
DUSt3R 3.73 97.8 7.30 91.6 4.96 96.4 3.94 96.6 2.55 99.6 25.4 64.2 16.9 76.2 6.68 92.6 8.93 89.4 6.62
Metric3D V2 3.94 97.6 3.50 98.4 3.24 99.0 3.28 98.3 2.10 99.4 26.6 71.7 7.15 94.8 2.75 98.7 6.57 94.7 3.64
UniDepth V1 3.40 98.6 3.55 98.7 4.92 97.5 3.76 98.2 2.48 99.9 24.9 64.1 9.46 90.8 4.90 96.2 7.17 93.0 3.62
Marigold 4.63 97.3 7.29 93.8 6.08 96.3 4.35 97.2 2.78 99.9 21.2 75.0 14.6 80.5 6.34 94.3 8.41 91.8 5.69
Geowizard 4.69 97.4 8.14 92.5 6.90 93.9 4.50 97.1 2.00 99.9 17.8 76.2 16.5 75.7 7.03 92.7 8.45 90.7 6.44
DA V1 3.82 98.3 5.04 96.4 6.23 95.2 4.23 97.3 1.98 100 20.1 71.8 11.3 86.1 6.75 92.6 7.43 92.2 4.83
  -metric indoor 3.82 98.3 9.95 86.5 6.23 95.2 4.23 97.3 1.98 100 20.1 71.8 17.0 74.0 6.75 92.6 8.76 89.5 -
  -metric outdoor 7.68 93.8 5.04 96.4 6.21 96.6 7.00 94.2 2.77 99.8 20.6 70.0 11.3 86.1 7.03 93.2 8.45 91.3 -
DA V2 4.16 97.9 6.77 94.3 4.63 97.2 3.44 98.3 1.44 100 17.1 76.6 13.4 81.8 5.41 94.6 7.04 92.6 -
  -metric indoor 4.16 97.9 7.09 92.3 4.63 97.2 3.44 98.3 1.44 100 17.1 76.6 14.3 79.8 5.41 94.6 7.20 92.1 -
  -metric outdoor 8.65 91.0 6.77 94.3 7.24 93.5 6.80 93.5 2.29 100 22.4 67.1 13.4 81.8 8.19 90.7 9.47 89.0 2.94
Ours 2.92 98.6 3.94 98.0 2.69 99.2 2.74 97.9 0.94 100 13.0 83.2 8.40 92.1 3.16 97.5 4.72 95.8 1.56